### v1

# External Attention Transformer: A Robust AI Model for Identifying Initial Eccentricity Signatures in Binary Black Hole Events in Simulated Advanced LIGO Data

Abstract
Initial orbital eccentricities of gravitational wave (GW) events associated with merging binary black holes (BBHs) should provide clues to their formation scenarios, mainly because various BBH formation channels predict distinct eccentricity distributions. However, searching for inspiral GWs from eccentric BBHs is computationally challenging due to sophisticated approaches to model such GW events. This ensures that Bayesian parameter estimation methods to characterize such events are computationally daunting. These considerations influenced us to propose a novel approach to identify and characterize eccentric BBH events in the LIGO-Virgo-KAGRA (LVK) collaboration data sets that leverages external attention transformer models. Employing simulated data that mimic LIGO O4 run, eccentric inspiral events modeled by an effective-one-body numerical- relativity waveform family, we show the effectiveness of our approach. By integrating this transformer-based framework with a convolutional neural network (CNN) architecture, we provide efficient way to identify eccentric BBH GW events and accurately characterize their source properties.
1 Introduction
A decade long routine detections of transient gravitational waves (GWs) from merging stellar mass binary black holes (BBHs) by the LIGO-Virgo-KAGRA (LVK) collaboration beginning with GW150914 have established the era of GW Astronomy [1]. This collaboration have so far detected and characterised approximately 100 gravitational wave (GW) events during the three concluded observing runs (i.e. O1-O3) [2, 3, 4]. The ongoing and planned observational runs are expected to should astrophysical evidences for the dominant formation channel for these BBHs[5]. We note that the LVK consortium is currently in the second half of the fourth observing run, and the detected gravitational wave (GW) candidates are listed on the GraceDB website  [^1] https://gracedb.ligo.org/superevents/public/O4/. The majority of these signals are most likely produced by binary black hole (BBH) mergers, with the black holes inspiraling in quasi-circular orbits.
It turns out that initial orbital eccentricities, component spins, masses, and eccentricities of compact binaries provide crucial clues for distinguishing between different formation channels, as they encapsulate information about the formation history of the system [6, 7]. As the eccentricity of a binary system rapidly decreases due to the emission of gravitational wave radiation [8], isolated binary black holes that inspiral in eccentric orbits are expected to have negligible eccentricities by the time they reach the sensitivity range of LVK observations. GW emissions from binaries formed through the isolated evolution process are expected to be detected as they inspiral on quasi-circular orbits. In this scenario, two stars evolve around each other with minimal interaction with their environment, eventually forming compact objects and merging within the age of the Universe [9, 6, 7]. However, in the case of dynamical formation, certain processes can cause the detected eccentricity to be non-negligible. Dynamical formation occurs in populated environments, such as globular clusters, young star clusters, and galactic nuclei [10, 11, 12]. Merger progenitors formed through the dynamical formation channel may be observed with eccentricities $e_{10}$ $>$ 0.1 in the LVK sensitive frequency range (above 10 Hz) [13, 14, 15, 16, 6, 17].
Interestingly, several detected events, such as GW190521 [18], have been argued to favor non-negligible orbital eccentricities within the LVK sensitive band (above 10 Hz)[19, 20, 21, 22, 23, 24]. In contrast, Ref. [25] re-analyzed 57 GW events using from O1 to O3 observing run of LVK collaboration and argued that three BBH events should have non-negligible eccentricities at 10Hz. Interestingly, these authors find no evidence for orbital eccentricity in GW190521. Further, it was argued recently that a BH-neutron star merger event, namely GW200105, should have a residual orbital eccentricity of $\sim 0.15$ around 20 Hz [26].
We note that the LVK collaboration typically employs matched-filtering techniques to detect and characterise transient GW events and the underlying template families involve compact binaries in quasi-circular orbits [4]. Further, parameter estimation pipelines, such as PyCBC [27] and Bilby [28], commonly employ Bayesian inference techniques to thoroughly explore the source parameter space of coalescing compact binaries consisting of BHs and neutron stars. However, the inclusion of eccentricity as an additional parameter substantially increases the computational cost. Ref. [29] shows that using quasi-circular template banks in search pipelines may lead to missing up to 2.2% of events, those with a log-uniform distribution of eccentricities. Additionally, this approach may introduce biases in the recovered parameters. As pointed out in Ref. [30], pipelines using quasi-circular GW templates are capable of detecting eccentric compact binaries with an eccentricity $e_{10}<0.1$ at a frequency of 10 Hz. For systems with larger eccentricities, it is necessary to conduct eccentricity-targeted searches by constructing templates that account for the effects of eccentricity on the waveform [31, 32, 33]. Rapidly identifying and characterising these eccentric mergers is therefore crucial for population studies and time-critical multimessenger follow-up. In this work we introduce an External-Attention Transformer that sidesteps expensive likelihood evaluations and provides real-time estimates of intrinsic BBH parameters such as initial eccentricity and source-frame chirp mass directly from GW spectrogram data.
The remainder of this paper is structured as follows. Section 2 details the model architecture and training procedure. Section  3 shows the results of our trained network and assesses its performance on realistic data. We conclude in Section  4.
2 Methods
Machine learning has been applied to GW data analysis. Some models demonstrate that machine learning can significantly reduce computational costs while maintaining sensitivity comparable to traditional GW data analysis methods. These techniques have been utilized in various tasks, including signal detection [34, 35, 36], glitch classification [37], parameter estimation [38, 39, 40] , and a range of other applications (see [41] and references therein).
2.1 Neural Network Architecture
Transformers [42] were originally developed as building blocks for natural language processing (NLP) tasks (e.g., GPT-4 [43]) and subsequently extended to fields such as computer vision [44] and time-series forecasting [45]. Recently, transformer-based models have been utilized in astronomy and cosmology (e.g. [46, 47, 48]). Although transformers are not yet as commonly used in these fields as other deep learning architectures like convolutional neural networks (CNNs) (e.g. in GW field [41] and references therein), they offer distinct advantages over recurrent neural networks (RNNs) such as long short-term memory (LSTM) [49] networks. Notably, transformers excel at capturing long-range dependencies and enable efficient parallel computation, whereas LSTMs inherently rely on sequential data processing.
The attention mechanism [42] is a fundamental component of transformer models, designed to capture relationships among different data segments or patches. This mechanism consists of queries, keys, and values. Queries act as prompts seeking specific information, keys function as reference indicators to pinpoint relevant responses, and values hold the underlying data to be accessed and processed. During training, the model learns to compare queries against keys to assess the relevance of each value, thereby enabling effective extraction and representation of meaningful relationships within the data. The self-attention matrix is computed using the softmax function of the scaled dot-product between query and key matrices, subsequently multiplied by the value matrix [42]:
$\mathrm{Attention_{self}}=\mathrm{softmax}\Big{(}QK^{T}\Big{)}V,$ (2.1)
Multi-head attention enhances the capability of single-head attention by projecting queries, keys, and values linearly through different learned projection weights, allowing multiple representations to be captured simultaneously during training. Self-attention mechanisms capture internal relationships within each sample individually, they inherently disregard relationships between different samples. To address this limitation, the external attention mechanism has been introduced by EANet [50], defined as [51]:
$\mathrm{Attention_{external}}=\mathrm{Norm}\Big{(}QM_{k}^{T}\Big{)}M_{v},$ (2.2)
where $M_{k}$ and $M_{v}$ are memory units representing key and value matrices, respectively. These parameters are learned independently from samples, enabling external attention to capture relationships across multiple samples and thereby enhancing generalization.
In this work, we employ an external transformer architecture adapted from the AENet classification model  [^2] https://keras.io/examples/vision/eanet/, modifying it specifically for our regression task. Additionally, we use two convolutional layers with 64 and 128 filters, respectively, each with a kernel size of $3\times 3$. Each convolutional layer is subsequently followed by batch normalization and max pooling. After these layers, we apply global average pooling and a dense output layer for the regression task. We also checked fully connected (dense) neural network layers as an alternative to the CNN, but found that the CNN architecture yielded better performance for our task.
Our goal is to evaluate whether external transformers can effectively identify characteristic patterns associated with orbital eccentricity in BBH GW events. It may be noted that temporally evolving GW polarization states associated with non-spinning BBHs merging along eccentric orbits show features that arise from orbital, periastron advance, and GW emission effects as explained in [52]. We generate GW spectrogram samples to create a dataset for training the model and assessing its performance. We then evaluate our model using new unseen mock GW data.
2.2 Preprocessing: creating the Synthetic Data
We employ the SEOBNRv5EHM waveform model [53, 54] from the effective-one-body numerical-relativity (EOBNR) waveform model family [55] using the pyseobnr package [56] and LALSuite [57, 58] to generate simulated eccentric gravitational wave (GW) waveforms. This model represents an order of improvement over the SEOBNRv4EHM model described in Ref. ([59], [60]) and includes higher-order multipoles: $(\ell,|m|)=(2,2),(2,1),(3,2),(3,3),(4,3),(4,4)$. For this work, we focus on the dominant $(\ell,|m|)=(2,2)$ mode in our simulations. The chirp mass $M_{c}$ of each binary is drawn uniformly from the range (15, 60) $M_{\odot}$. The mass ratio $q=m_{2}/m_{1}$ spans the range 0.25 to 1. The initial orbital phase $\zeta$ is sampled uniformly over $(0,2\pi)$, and the eccentricity at 15 Hz, $e_{15,\mathrm{Hz}}$, is drawn uniformly from $(0.001,0.35)$. Distances are randomly sampled within the range 100 to 1500 Mpc. The sky position (right ascension and declination) and orientation (inclination and polarization angles) are each uniformly sampled over the sky and the polarization sphere, respectively. For simplicity, we neglect the spins of the binary black hole components. Waveforms are generated at a sample rate of 4096 Hz, with a frequency resolution of 1/16 Hz. Following the methodology outlined in Ref. [61], we generate spectrograms of the simulated signals using PyCBC [62]. Each waveform is 30 seconds long, spanning the time window $[-15,15]$ seconds, and is injected into 40 seconds of Gaussian noise ($[-20,20]$ seconds) using the O4 sensitivity curve for the LIGO detectors [^3] https://dcc.ligo.org/ligo-t2000012/public. The merger time is set to zero for all synthetic signals. To reduce computational costs, we consider only the data in the $[-10,2]$ seconds window for generating spectrograms. We have checked that this time range is sufficient to capture the full signal duration, even for the longest signals in our simulations, which correspond to $M_{c}$=15 $M_{\odot}$ and q=1. For each simulated dataset, we calculate the matched-filter signal-to-noise ratio (SNR) and retain only those samples with LIGO Hanford–LIGO Livingston network SNR greater than 10. The resulting signals are band-pass filtered in the $[20,512]$ Hz frequency range and whitened, after which the Q-transform is applied to the whitened signal. The resulting Q-transform spectrograms are resized to $256\times 256$ pixels and stacked to construct the dataset for subsequent visual processing tasks.
To construct our dataset for training and testing the model, we repeat the above simulation procedure to generate 5000 events for each network SNR range: (10–15), (15–20), (20–30), (30–40), (40–50), and (50–60). This results in a total of 30,000 samples, with each sample represented as a spectrogram of shape (256, 256, 2). The three dimensions correspond to frequency, time, and the two LIGO detectors (Hanford and Livingston), respectively.
2.3 Training
We divided the simulated GW spectrogram dataset into three subsets: training, testing, and validation. We randomly allocated 80% of the total sample for training and 20% for testing. Additionally, 20% of the training data was set for validation. To ensure consistent input distributions, we apply z-score normalization to all data subsets: we first compute the mean and standard deviation on the training set, then subtract the mean and divide by the standard deviation for every sample in the training, validation, and test sets. Our model is designed for regression, with a primary focus on eccentricity estimation. However, as shown in Ref. [63, 64], there exists a degeneracy between chirp mass and eccentricity. To address this, we include both eccentricity and chirp mass as target variables in our model, and both parameters are standardized using the z-score normalization method. During inference, we recover the physical values of the predicted parameters by applying the inverse transformation of the z-score normalization.
We fine-tune several hyperparameters of the AENet model  [^4] https://keras.io/examples/vision/eanet/ to optimize its performance for our specific objectives. We set the maximum training epochs to 500, implementing early stopping if no improvement in validation loss occurs over 15 consecutive epochs.
During training, we utilize the Adam optimizer with a learning rate of 0.001, a weight decay of 0.0001, and a batch size of 32. We further configure the training process to reduce the learning rate by a factor of two whenever the validation loss plateaus for 10 epochs. Our network divides each spectrogram into smaller patches, each sized $16\times 16$ pixels. These patches then pass through an embedding layer with a projection dimension of 128, followed by multi-head external attention with eight heads.
The mean squared error (MSE) loss function used for our regression task is defined as:
$LOSS_{(MSE)}=\frac{1}{N_{\rm batch}}\sum_{j=1}^{N_{\rm batch}}\sum_{i=1}^{2}\left(y_{i,j}^{\rm pred}-y_{i,j}^{\rm truth}\right)^{2}~{},$ (2.3)
where $y_{1}$ and $y_{2}$ represent the eccentricity and chirp mass, respectively, and $N_{\mathrm{batch}}$ is the number of data in each batch.
The model training is performed on a single GPU (e.g., NVIDIA GeForce RTX 2070 with 8 GB VRAM) utilizing 80% of the dataset (approximately 24,000 gravitational wave spectrograms) with an additional 20% of this subset used for validation. The entire training procedure takes less than one hour. The model has $1,840,450$ trainable parameters.
Figure 1 shows the training and validation loss curves as a function of epoch. During training, we monitor the validation loss and select the final model corresponding to the lowest minimum of the validation loss. Figure 2 presents scatter plots comparing the model-predicted eccentricity and chirp mass with the true injected values from the test data. The Advanced LIGO-Virgo-KAGRA (LVK) detectors are sensitive to orbital eccentricities greater than 0.05 at a frequency of 10 Hz [65]. Therefore, we present results only for eccentricities above this threshold.
We perform inference with our trained model for $e_{15,\mathrm{Hz}}$ values in the range [0.05, 0.32] and $M_{c}$ values in the range [15, 55] $M_{\odot}$. Although the model is trained on a broader range of these parameters, we restrict the inference range to avoid systematic bias arising from edge effects in the data [66]. In the right panel of Figure 2, we observe increased scatter in the recovered $M_{c}$ values at higher masses. This is expected, as the time-frequency maps become shorter for higher-$M_{c}$ binary black holes, leading to greater uncertainties in the recovered $M_{c}$ values with our model. In the left panel of Figure 2, we find that for low/high eccentricity, the model tends to overestimate/underestimate values when the network SNR is low, indicating difficulty in extracting reliable eccentricity information from low-quality data. Therefore, we need to investigate to determine if there are any significant systematic biases in the estimation. In the next section, we examine the distributions of eccentricity and chirp mass in greater detail.
Figure 1: The evolution of MSE loss function during training process. We save and use the model at the epoch with lowest validation loss. The blue line shows the train loss and the orange line represents the validation loss. The red dashed line marks the epoch where the validation loss reaches its lowest value.
Figure 2: A comparison between the input (injected BBH) eccentricity and chirp mass and the corresponding values predicted by the model from the test datasets. Blue points represent injection samples with 10 < SNR < 20 and brown points represent samples with 20 < SNR < 60.
3 Implications of Our Approach
To evaluate the performance of our model, we generate several independent test datasets, each with fixed values of chirp mass and eccentricity, across different network SNR ranges: (10–15), (15–20), (20–30), (30–40), (40–50), and (50–60). For each network SNR range, we generate 500 events. On average, the model required on the order of 100 ms to process a single event, as determined from 100 repeated inferences. Figure 3 shows the recovered distributions of $e_{15,\mathrm{Hz}}$ = 0.07 and $M_{c}$ = 40 $M_{\odot}$ for 500 events in each network SNR bin. In each subpanel, the parameters are recovered within 1 $\sigma$, and the size of the contours decreases as the network SNR increases, indicating improved precision. Figure 4 presents a similar analysis for a higher eccentricity value of 0.2.
[…]
4 Conclusions
Most LVK pipeline searches currently neglect the eccentricity parameter in gravitational wave (GW) signals due to the significant computational overhead associated with incorporating this parameter into Bayesian inference methods for parameter estimation. Additionally, waveform models accounting for eccentricity are computationally intensive and relatively slow. Consequently, standard analysis pipelines typically assume quasi-circular orbits for parameter estimation. However, neglecting eccentricity in the parameter estimation of genuinely eccentric binary black hole (BBH) events can introduce biases in the recovered parameters [29].
In this study, we utilize an external transformer-based neural network to effectively identify the eccentricity of BBH systems from GW spectrogram data. Our approach modifies the external transformer classification model, initially introduced by Ref. [50], converting it into a regression model with the addition of convolutional neural network (CNN) layers. This enhanced model successfully estimates both chirp mass and eccentricity directly from GW spectrograms.
A significant advantage of our method is its rapid and computationally efficient training process, enabled by the external attention mechanism, which employs external memory to capture general patterns effectively across subsamples. Specifically, we trained the model on 24,000 gravitational wave spectrograms, corresponding to the training dataset, using a single GPU (NVIDIA GeForce RTX 2070 with 8 GB VRAM). During training, 20% of the training data was used for validation, and the entire training process was completed in less than one hour. The average inference time per event, measured over 100 runs, was on the order of 100 ms. Testing across diverse datasets, as discussed in Section 3, demonstrates that our model achieves sufficient precision in determining eccentricity. Recognizing the intrinsic degeneracy between eccentricity and chirp mass, our approach simultaneously estimates both parameters to improve robustness and reliability. Furthermore, the model is trained on data simulating the current LVK O4 observational run, making it immediately applicable to ongoing analyses. Due to its efficient training capability, this model can easily be retrained or updated for future LVK observational runs, highlighting its practical utility for rapid and accurate GW parameter estimation. These considerations imply that our approach should allow us to provide independent constraints on the reported initial eccentricities of a few O3 GW events [18, 25, 26]. This is currently under investigation.
[…]
Acknowledgments
E.K. and H.M.L. are supported by the National Research Foundation of Korea 2021M3F7A1082056. C.G.S. acknowledge support from the Basic Science Research Program (2018R1A6A1A06024977 and RS-2025-00515276, respectively) through Korea’s NRF funded by the Ministry of Education.

### v2

@@
 Dynamical formation occurs in populated environments, such as globular clusters, young star clusters, and galactic nuclei [10, 11, 12].
-Merger progenitors formed through the dynamical formation channel may be observed with eccentricities $e_{10}$ $>$ 0.1 in the LVK sensitive frequency range (above 10 Hz) [13, 14, 15, 16, 6, 17].
+Merger progenitors formed through the dynamical formation channel may be observed with eccentricities $e_{10}$ >> 0.1 in the LVK sensitive frequency range (above 10 Hz) [13, 14, 15, 16, 6, 17].
 Interestingly, several detected events, such as GW190521 [18], have been argued to favor non-negligible orbital eccentricities within the LVK sensitive band (above 10 Hz)[19, 20, 21, 22, 23, 24].
@@
 [61], we generate spectrograms of the simulated signals using PyCBC [62].
-Each waveform is 30 seconds long, spanning the time window $[-15,15]$ seconds, and is injected into 40 seconds of Gaussian noise ($[-20,20]$ seconds) using the O4 sensitivity curve for the LIGO detectors [^3] https://dcc.ligo.org/ligo-t2000012/public.
+Each waveform is 30 seconds long, spanning the time window [15,15][-15,15] seconds, and is injected into 40 seconds of Gaussian noise ([20,20][-20,20] seconds) using the O4 sensitivity curve for the LIGO detectors [^3] https://dcc.ligo.org/ligo-t2000012/public.
 The merger time is set to zero for all synthetic signals.
-To reduce computational costs, we consider only the data in the $[-10,2]$ seconds window for generating spectrograms.
+To reduce computational costs, we consider only the data in the [10,2][-10,2] seconds window for generating spectrograms.
 We have checked that this time range is sufficient to capture the full signal duration, even for the longest signals in our simulations, which correspond to $M_{c}$=15 $M_{\odot}$ and q=1.
@@
 To evaluate the performance of our model, we generate several independent test datasets, each with fixed values of chirp mass and eccentricity, across different network SNR ranges: (10–15), (15–20), (20–30), (30–40), (40–50), and (50–60).
-For each network SNR range, we generate 500 events.
+For each network SNR range, we generate 300 events.
 On average, the model required on the order of 100 ms to process a single event, as determined from 100 repeated inferences.
-Figure 3 shows the recovered distributions of $e_{15,\mathrm{Hz}}$ = 0.07 and $M_{c}$ = 40 $M_{\odot}$ for 500 events in each network SNR bin.
+We trained our model independently 10 times, and for each run, we saved the best-performing model based on the minimum validation loss, following the procedure described in subsection 2.3.
+Figure 3 shows the recovered distributions of $e_{15,\mathrm{Hz}}$ = 0.07 and $M_{c}$ = 40 $M_{\odot}$ for 300 events in each network SNR bin.
 In each subpanel, the parameters are recovered within 1 $\sigma$, and the size of the contours decreases as the network SNR increases, indicating improved precision.
@@
 […]
+[…]
+To better characterize prediction confidence in Figures 3-6, we evaluate model uncertainty using a deep-ensemble approach.
+We performed inference using all 10 trained models on several datasets and computed the standard deviation of the predictions for each event.
+We find that when the SNR is low, the main issue is the strong noise in the data.
+This noise makes it difficult for the model to identify true patterns, leading to greater variation in its predictions.
+In other words, most of the uncertainty at low SNR is due to the data, not limitations of the model.
+As the SNR increases and the signal becomes more prominent relative to the noise, the model’s predictions become much more precise and confident.
+In this high SNR regime, model uncertainty is comparable to the total uncertainty.
+The ensemble shows no significant improvement over a single model.
+3.1 Quantitative Performance Metrics
+To further quantify model performance, we report the mean absolute error (MAE), root mean squared error (RMSE), and their associated 95% confidence intervals across bins of $e_{15,\mathrm{Hz}}$ and $M_{c}$, separately for low and high SNR categories.
+For this evaluation, we generated a new test dataset, with the data binned by $e_{15,\mathrm{Hz}}$, $M_{c}$ , and SNR using the following bin edges: $e_{15,\mathrm{Hz}}$ bins of (0.05–0.10), (0.10–0.15), (0.15–0.20), (0.20–0.25), and (0.25–0.30); $M_{c}$ bins of (15–25), (25–35), (35–45), and (45–55) $M_{\odot}$; and SNR bins of (10–15), (15–20), (20–30), (30–40), (40–50), and (50–60).
+We used $N=100$ test samples in each $(e_{15\,\mathrm{Hz}},\,M_{c},\,\mathrm{SNR})$ bin combination.
+Throughout this analysis, we define low SNR as $10<SNR<20$ and high SNR as $20<SNR<60$.
+The detailed results are summarized in Table 1 for $e_{15,\mathrm{Hz}}$ bins and Table 2 for $M_{c}$ bins.
+These quantitative metrics provide a rigorous assessment of the model’s predictive accuracy across the relevant parameter space.
+As expected, the model achieves better performance for lower $M_{c}$ systems, whereas the errors increase for higher $M_{c}$.
+This trend, discussed in Section 2.3, arises because the time-frequency maps become shorter for higher $M_{c}$, leading to greater uncertainties in the recovered $M_{c}$.
+For $e_{15,\mathrm{Hz}}$ the model shows consistently reliable performance across the explored range.
+Table 1: MAE and RMSE for $e_{15,\mathrm{Hz}}$ bins with 95% confidence intervals over ten independent runs.
+[…]
+0.05–0.10 $0490^{+0.0027}_{-0.0032} 0.0$$637^{+0.0033}_{-0.0029} 0.0$$216^{+0.0009}_{-0.0010} 0.0$$0281^{+0.0015}_{-0.0012}$
+0.10–0.15 $0424^{+0.0022}_{-0.0021} 0.0$$529^{+0.0025}_{-0.0026} 0.0$$244^{+0.0011}_{-0.0011} 0.0$$0316^{+0.0015}_{-0.0013}$
+0.15–0.20 $0422^{+0.0020}_{-0.0020} 0.0$$519^{+0.0022}_{-0.0024} 0.0$$247^{+0.0012}_{-0.0011} 0.0$$0323^{+0.0017}_{-0.0016}$
+0.20–0.25 $0421^{+0.0023}_{-0.0023} 0.0$$531^{+0.0026}_{-0.0025} 0.0$$206^{+0.0009}_{-0.0010} 0.0$$0265^{+0.0013}_{-0.0013}$
+0.25–0.30 $0415^{+0.0023}_{-0.0025} 0.0$$535^{+0.0027}_{-0.0032} 0.0$$181^{+0.0011}_{-0.0008} 0.0$$0251^{+0.0018}_{-0.0019}$
+Table 2: MAE and RMSE for $\mathcal{M}_{c}$ bins with 95% confidence intervals over ten independent runs.
+[…]
+15–25 $0.8002^{+0.0500}_{-0.0460}$ $1.1008^{+0.1333}_{-0.0995}$ $0.4813^{+0.0199}_{-0.0184}$ $0.6084^{+0.0208}_{-0.0210}$
+25–35 $1.3042^{+0.0871}_{-0.0776}$ $1.8262^{+0.1346}_{-0.1316}$ $0.6030^{+0.0257}_{-0.0268}$ $0.8116^{+0.0636}_{-0.0556}$
+35–45 $2.4309^{+0.1292}_{-0.1289}$ $3.2095^{+0.1688}_{-0.1703}$ $1.1561^{+0.0607}_{-0.0655}$ $1.6413^{+0.0921}_{-0.0912}$
+45–55 $2.8756^{+0.1432}_{-0.1326}$ $3.6514^{+0.1905}_{-0.1673}$ $1.6351^{+0.0773}_{-0.0662}$ $2.1609^{+0.1087}_{-0.0938}$
 4 Conclusions
@@
 The average inference time per event, measured over 100 runs, was on the order of 100 ms.
+We implemented ensemble learning to quantify model uncertainty.
 Testing across diverse datasets, as discussed in Section 3, demonstrates that our model achieves sufficient precision in determining eccentricity.
+Our findings show that the ensemble method does not improve performance compared to the single model.
 Recognizing the intrinsic degeneracy between eccentricity and chirp mass, our approach simultaneously estimates both parameters to improve robustness and reliability.